OpenStudy (anonymous):

f(x) = x^2 + 6x - 16 two real rational solutions g(x) = x^2 +6x + 1 two real irrationals solutions i found that f(x) is two real rational solutions and that g(x) is two real irrationals solutions on another open study question but dont understand how to get that

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

You can apply the quadratic formula: If you have a polynomial of the form \(ax^2+bx+c\), then \[ x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] For \(f(x) = x^2 + 6x - 16 \), \(a=1, b=6, c=-16\) \[ x=\frac{-6\pm\sqrt{6^2-4(1)(-16)}}{2(1)}\\ x=\frac{-6\pm\sqrt{100}}{2}\\ x=\frac{-6\pm 10}{2}\\ x=\frac{-6-10}{2}, or ~ x=\frac{-6+10}{2}\] For \(g(x) = x^2 +6x + 1\), \(a=1, b=6, c=1\) \[ x=\frac{-6\pm\sqrt{6^2-4(1)(1)}}{2(1)}\\ x=\frac{-6\pm\sqrt{32}}{2}\\ x=\frac{-6-\sqrt{32}}{2}, or ~ x=\frac{-6+\sqrt{32}}{2}\] But \(\sqrt{32}=4\sqrt{2}\) and is irrational because \(\sqrt{2}\) is irrational. So thus, the roots above will also be irrational.

OpenStudy (anonymous):

@kirbykirby how would i convert f(x) into general, vertex form

OpenStudy (kirbykirby):

This would involve the process of completing the square. For \( x^2 + 6x - 16\), first take the \(b\) coefficient (\(6\)), and add and subtract \(\left(\dfrac{b}{2}\right)^2\), so \(\left(\dfrac{6}{2}\right)^2=9\) Thus: \[ x^2+6x\color{red}{+9-9}-16\] Now the first 3 terms will give you a binomial of the form \((a+b)^2\). Since this is equal to \(a^2+2ab+b^2\), then it follows from the 1st three terms above that \(a=1\) and \(b=3\) \[ (x+3)^2-25\], the -25 comes from combining -9 and -16 So it is in vertex form \(a(x-h)^2+k\) where \(a=1, h=-3, k=-25\)